What is the probability of flipping a coin and getting heads, and then rolling a six-sided


die and getting a result with two or fewer pips showing


1
Expert's answer
2021-12-14T13:51:10-0500

the probability of flipping a coin and getting head:

"P=0.5"

probability of getting a result with two or fewer pips showing:

"P=\frac{1+2+1}{36}=1/9"


total probability:

"1/2\cdot 1/9=1/18"